# DiBS-MTL: Transformation-Invariant Multitask Learning with Direction Oracles

### v1

@@ 5 Experiments
We now evaluate DiBS-MTL in standard MTL benchmarks extensively used in literature (Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023), a demonstrative two objective example, a computer vision benchmark (NYU-v2) (Silberman et al., 2012), and a multitask reinforcement learning dataset (Meta-World MT10) (Yu et al., 2020b)  [^1] Code for all experiments is provided at https://github.com/suryakmurthy/dibs-mtl. Our main aims are: (i) to verify that DiBS-MTL  reliably converges to Pareto solutions from varying initializations, (ii) to compare the overall performance of solutions identified by DiBS-MTL  to those produced by state of the art approaches across a variety of large-scale MTL benchmarks, and (iii) to verify the robustness of DiBS-MTL  to nonaffine task loss transformations that skew the results of other methods.
Figure 1: The loss functions used in the demonstrative two-dimensional example. The transformation we use is $h(\mathcal{L})=\operatorname{sign}(\mathcal{L})\cdot\mathcal{L}^{4}$.
(a) Original loss pair $(\mathcal{L}_{1},\mathcal{L}_{2})$
(b) Transformed loss pair $(h(\mathcal{L}_{1}),\mathcal{L}_{2})$
Figure 2: Results for the demonstrative two objective nonconvex problem. Each initialization is represented by $\bullet$, and    denotes the Pareto front. We retain the original $\mathcal{L}_{1}$ axis in the transformed case to better visualize the different results reached by baselines. DiBS-MTL provides balanced Pareto solutions invariant to monotonic nonaffine task-loss transformations, while competing methods display degraded fairness, favoring task 1.
5.1 Does DiBS-MTL reliably converge to Pareto solutions?
@@
We first test DiBS-MTL on a demonstrative two-dimensional, nonconvex multi-objective optimization example with two objectives $(\mathcal{L}_{1},\mathcal{L}_{2})$, which is a standard benchmark for illustrating the ability of MTL methods to achieve balanced Pareto solutions (Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023). The explicit forms of the objectives are provided in Appendix D.1. As shown in Figure 1, each objective in this example has deep valleys, with the bottoms having a large magnitude difference between the objective gradients, and a high (positive) curvature. Such phenomenon has been documented to exist when optimizing neural networks as well, and can lead to one objective (task) dominating others (Goodfellow et al., 2014; Yu et al., 2020a).
@@
We test DiBS-MTL along with 5 baselines—(1) linear scalarization (LS), which optimizes the unweighted loss average, (2) CAGrad (Liu et al., 2021a), (3) PCGrad (Yu et al., 2020a), (4) Nash-MTL (Navon et al., 2022), and (5) FAMO [^2] We were unable to reproduce the baseline plots reported in the original FAMO paper (Liu et al., 2023). The results shown here were obtained by strictly following the default parameters and instructions in the publicly available FAMO repository https://github.com/Cranial-XIX/FAMO. (Liu et al., 2023).
@@
DiBS-MTL yields Pareto solutions from diverse initializations, which are unaffected by monotonic nonaffine transformations.
@@
Figure 2 plots the solutions generated by DiBS-MTL for different initializations, for the costs $(\mathcal{L}_{1},\mathcal{L}_{2})$ above, as well as for $\left(h\left(\mathcal{L}_{1}\right),\mathcal{L}_{2}\right)$ where we apply the monotonic nonaffine transformation $h(x)=\operatorname{sign}(x)\cdot x^{4}$ to $\mathcal{L}_{1}$. We observe that DiBS-MTL consistently leads to balanced Pareto solutions for all initializations, and these solutions remain invariant to the transformation. It is also observed that unlike DiBS-MTL, all other MTL methods produce significantly biased solutions once $\mathcal{L}_{1}$ is transformed (i.e., comparing the top and bottom rows of Figure 2), signaling that they are susceptible to task domination when tasks are differently scaled.
Table 1: Results for supervised learning NYU-v2 benchmark. DiBS-MTL is competitive across all tasks, and displays healthy overall performance, with the best average relative performance drop ($\Delta m\%$), and second-best mean rank (MR). Bold, underlined values indicate best, second-best performances per column respectively.
@@
STL 38.30 63.76 0.6754 0.2780 19.21 25.01 69.15 57.20 30.14
MGDA 32.03 60.77 0.6103 0.2453 19.00 24.64 69.83 57.78 30.55 -0.6944 3.22
UW 39.08 64.73 0.5464 0.2285 23.04 27.34 62.85 49.23 23.49 3.7667 4.89
NashMTL 40.16 65.65 0.5332 0.2204 19.96 25.25 68.29 55.72 28.62 -3.9833 2.33
FAMO 37.58 64.08 0.5595 0.2297 19.15 25.04 69.36 57.44 30.23 -3.8200 3.33
LS 40.16 65.63 0.5446 0.2223 23.03 27.50 62.66 49.32 23.61 3.0569 4.22
DiBS-MTL 40.92 66.60 0.5337 0.2217 20.06 25.35 68.15 55.54 28.54 -4.1140 2.89
5.2 How does DiBS-MTL compare to existing MTL methods?
@@
To compare DiBS-MTL with existing methods in large-scale learning examples, we perform experiments in supervised learning (computer vision) and multitask reinforcement learning (MTRL) domains that are standard benchmarks in MTL literature (Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023). We begin by describing the set-up, baselines and metrics for both. We keep the default training procedures established in prior work for a fair comparison. All implementation details are given in Appendix D.
Supervised learning set-up.
The setting is a supervised learning computer vision problem with 3 tasks. Specifically, we use the NYU-v2 indoor scene dataset (Silberman et al., 2012) with 1449 RGBD images and dense per-pixel labeling for 3 learning tasks—semantic segmentation (13 classes), depth estimation, and surface normal prediction. Similar to prior work, we train a multitask attention network (MTAN) (Liu et al., 2019).
Supervised learning baselines.
We use 5 baselines—(1) LS, (2) uncertainty weighting (UW) (Kendall et al., 2018), and 3 methods which have consistently outperformed other MTL approaches for this particular dataset in prior work—(3) MGDA (Sener & Koltun, 2018), (4) Nash-MTL (Navon et al., 2022), and (5) FAMO (Liu et al., 2023).
Supervised learning metrics.
We report several task-specific metrics covering each prediction type. For semantic segmentation, we use the mean Intersection-over-Union (mIoU), which averages the per-class overlap between predicted and ground-truth regions, and the overall pixel accuracy across the image. For depth estimation, we report the mean absolute error and mean relative error. For surface normal prediction, we include the mean and median angular error between estimated and true normals, along with the percentage of pixels whose angular error falls below $30^{\circ}$, $22.5^{\circ}$, and $11.25^{\circ}$ (Silberman et al., 2012). In addition, we report two MTL-specific metrics used in previous work—average relative per-task performance drop $\Delta m\%$, and mean rank (MR) across tasks (Navon et al., 2022; Liu et al., 2023). $\Delta m\%$ is calculated relative to STL—single-task learning, corresponding to learning a separate model for each task—given for a method $m$ as $\Delta m\%=\left(\nicefrac{{1}}{{N}}\cdot\sum_{k\in[N]}\nicefrac{{\left(M_{m,k}-M_{\texttt{STL},k}\right)}}{{M_{\texttt{STL},k}}}\right)\times 100$, where $M_{i,k}$ is method $i$’s value on metric $M_{k}$. All reported results have been averaged over 3 seeds, as done in prior work.
MTRL set-up.
The setting is a multitask reinforcement learning (MTRL) problem with 10 tasks. Following prior work, we evaluate DiBS-MTL on the Meta-World MT10 benchmark (Yu et al., 2020b), where a robot arm manipulator has 10 tasks, each with distinct reward functions. The complete list of tasks is given in Appendix D.
MTRL metric and baselines.
We compare DiBS-MTL with (1) LS, (2) UW, and (3) FAMO. We note that results for Nash-MTL could not be collected due to its excessive run-time for this MTRL example (c.f., Section E.1 for a runtime comparison, where Nash-MTL is slow even for the simpler demonstrative example). Following prior work, for every method, we use Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as the underlying reinforcement learning algorithm. Following prior work, we report the fraction of tasks successfully completed  (Navon et al., 2022; Liu et al., 2023), where the definition of task success is as described in the Meta-World benchmark (Yu et al., 2020b). All reported results have been averaged over 10 random seeds. The implementation details are included in Appendix D.
DiBS-MTL performs competitively to state-of-the-art MTL methods.
Table 1 shows the results for the 3 task supervised learning NYU-v2 dataset. We find that DiBS-MTL performs consistently well for all tasks and is competitive with the baselines. Notably, DiBS-MTL achieves state-of-the-art performance in $\Delta m\%$ (overall average relative per-task performance drop), and second-best MR (mean rank) overall—both these metrics correspond to how well rounded an MTL method is with respect to all tasks. DiBS-MTL also achieves the state-of-the-art, and second-best performance in the segmentation and depth estimation tasks, respectively. In the 10 task robotic MTRL experiment as well, Table 2 and Figure 3(a) show that DiBS-MTL performs competitively, and achieves the second-best performance for the nominal case, where the reward functions for all tasks are the ones provided in Meta-World.
DiBS-MTL (ours)    FAMO    UW    LS
(a) Nominal (original rewards)
(b) Transformed (reach)
(c) Transformed (peg insert side)
(d) Transformed (window open)
Figure 3: Training Curves for MTRL Meta-World MT10 benchmark. Reward transformations significantly degrade performance of baseline MTL methods, while DiBS-MTL displays robustness.
5.3 How robust are DiBS-MTL & existing MTL methods against nonaffine transformations?
We now conduct experiments in the MTRL Meta-World MT10 benchmark Yu et al. (2020b) to investigate how DiBS-MTL and baseline MTL methods perform when some tasks undergo nonaffine transformations. We begin by describing the setup for the transformations. All training procedures are taken to be the same as in Section 5.2. We limit our study in this section to MTRL, because task-specific losses in computer vision are relatively standardized in the literature and thus transformations there will be unwarranted (Wang et al., 2022; Azad et al., 2023). In comparison, as we mention in Section 1, in the MTRL setting, rewards for different tasks might be designed on very different scales.
Transformed MTRL setup.
Reward transformations.
We consider the following task-transformation pairs: (i) $h(r)=\operatorname{sign}(r)\cdot r^{4}$ for reach, (ii) $h(r)=(5+r)^{4}$ for peg insert side, ($h$ is monotonic over the reward range), and (iii) $h(r)=\exp(r)$ for window open. These transformations were arbitrarily selected to illustrate a broad range of functional forms and tasks. We use the same baselines, metric, seed values, and training procedure as the nominal MTRL experiment in Section 5.2.
Table 2: Best-checkpoint success (evaluated every 200 episodes) results for MTRL Meta-World MT10 benchmark. Results are mean $\pm$ standard error over 10 seeds. Bold values indicate best performances per column.
Method Nominal Reach Window Open Peg Insert
DiBS-MTL $0.890\pm 0.023$ $\bm{0.850\pm 0.022}$ $\bm{0.920\pm 0.025}$ $\bm{0.700\pm 0.037}$
FAMO $\bm{0.920\pm 0.020}$ $0.700\pm 0.075$ $0.830\pm 0.033$ $0.570\pm 0.070$
LS $0.850\pm 0.040$ $0.480\pm 0.044$ $0.200\pm 0.026$ $0.230\pm 0.030$
UW $0.830\pm 0.033$ $0.650\pm 0.081$  [^3] UW results could not be collected for the window open case, because the training iterations became too unstable and the underlying MuJoCo simulator (Todorov et al., 2012) consistently gave out-of-bounds errors. $0.420\pm 0.074$
DiBS-MTL retains performance under nonaffine reward transformations.
Figures 3(a), 3(b), 3(c) and 3(d) and Table 2 show that DiBS-MTL achieves remarkable robustness to the different reward transformations compared to existing MTL methods. DiBS-MTL achieves state-of-the-art overall performance for all transformed cases, while other methods face significant performance degradation in one or more cases compared to their performance in the nominal setting.

### v2

@@ 5 Experiments
We now evaluate DiBS-MTL in standard MTL benchmarks extensively used in literature (Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023), a demonstrative two objective example, a multi-task reinforcement learning benchmark (Meta-World MT10) (Yu et al., 2020b), and supervised learning computer vision and graph neural network benchmarks (NYU-v2, Cityscapes and QM9) (Silberman et al., 2012; Cordts et al., 2016; Blum and Reymond, 2009). The results for the Cityscapes benchmark are included in Appendix D. Our main aims are: (i) to verify that DiBS-MTL  consistently converges to Pareto-optimal solutions from diverse initializations in nonconvex settings, (ii) to show that due to robustness against monotonic nonaffine task loss transformations, DiBS-MTL significantly outperforms state-of-the-art MTL baselines in examples where task losses are poorly scaled relative to each other, and (iii) to confirm that leveraging normalized gradients to achieve this robustness does not compromise the solution quality of DiBS-MTL  and that DiBS-MTL is competitive to state-of-the-art baselines across large-scale MTL benchmarks.
Figure 1: Loss functions in the nonconvex example (Section 5.1). We employ the nonaffine transformation $h(\mathcal{L}_{1})=\operatorname{sign}(\mathcal{L}_{1})\cdot\mathcal{L}_{1}^{4}$.
Figure 2: DiBS-MTL successfully avoids task domination due to its invariance to monotonic nonaffine transformations, while baseline MTL methods visibly favor task 1 when $\mathcal{L}_{1}$ is transformed, and some baseline methods even fail to reach the Pareto front. $\bullet$ and    denote initializations and the Pareto front respectively. To better visualize results for the transformed case, we retain the original $\mathcal{L}_{1}$ axis. Results for multi-step DiBS-MTL and additional baselines are included in Figure 6.
5.1 Does DiBS-MTL consistently converge to Pareto solutions from diverse initializations?
@@
We first test DiBS-MTL on a challenging two-dimensional, nonconvex multi-objective optimization example with two objectives $(\mathcal{L}_{1},\mathcal{L}_{2})$, which is a standard benchmark for illustrating the ability of MTL methods to achieve balanced Pareto solutions (Yu et al., 2020a; Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023; Ban and Ji, 2024). As shown in Figure 1, each objective (provided in Appendix E.1) in this example has deep valleys, with the bottoms having a large magnitude difference between the objective gradients, and a high (positive) curvature. Such phenomenon is documented to exist when optimizing neural networks as well, and can lead to one objective (task) dominating others (Goodfellow et al., 2014; Yu et al., 2020a).
@@
We test 5-step DiBS-MTL along with 5 baselines—(1) linear scalarization (LS), which optimizes the unweighted loss average, (2) CAGrad (Liu et al., 2021a), (3) FairGrad (Ban and Ji, 2024), (4) Nash-MTL (Navon et al., 2022), and (5) FAMO [^1] We were unable to reproduce the baseline plots reported in the original FAMO paper (Liu et al., 2023). The results shown here were obtained by strictly following the default parameters and instructions in the publicly available FAMO repository https://github.com/Cranial-XIX/FAMO. (Liu et al., 2023).
@@
DiBS-MTL yields Pareto solutions from diverse initializations, and unlike other methods, DiBS-MTL is invariant to monotonic nonaffine objective transformations.
@@
Figure 2 plots the solutions generated by DiBS-MTL for different initializations, for the costs $(\mathcal{L}_{1},\mathcal{L}_{2})$ above, as well as for $\left(h\left(\mathcal{L}_{1}\right),\mathcal{L}_{2}\right)$ where we apply the monotonic nonaffine transformation $h(x)=\operatorname{sign}(x)\cdot x^{4}$ to $\mathcal{L}_{1}$. We observe that DiBS-MTL consistently leads to balanced Pareto solutions for all initializations, and these solutions remain invariant to the transformation. It is also evident that unlike DiBS-MTL, all MTL baselines produce significantly biased solutions once $\mathcal{L}_{1}$ is transformed (i.e., comparing the top and bottom rows of Figure 2), signaling that baselines are susceptible to task domination when tasks are differently scaled. Further, DiBS-MTL achieves faster runtimes than the previous bargaining-inspired MTL method Nash-MTL (see Section F.1).
Table 1: 1-step DiBS-MTL achieves robustness to the different reward transformations on the MT10 multi-task reinforcement learning benchmark, and significantly outperforms all MTL baselines for all transformed settings. We report best-checkpoint success (evaluated every 200 episodes) mean results $\pm$ standard error over 10 seeds. Bold values indicate best performances per column.
For certain transformations, training with UW, Nash-MTL, FairGrad, and CAGrad was highly sensitive to loss scaling. This instability caused the training process to diverge, leading the underlying MuJoCo simulator (Todorov et al., 2012) to repeatedly produce out-of-bounds errors. As a result, we were unable to collect results for these cases.
Transformed task
No task transformed reach transformed window open transformed peg insert transformed
FAMO $0.920\pm 0.020$ $0.700\pm 0.075$ $0.830\pm 0.033$ $0.570\pm 0.070$
LS $0.850\pm 0.040$ $0.480\pm 0.044$ $0.200\pm 0.026$ $0.230\pm 0.030$
UW $0.830\pm 0.033$ $0.650\pm 0.081$ $0.420\pm 0.074$
FairGrad $0.950\pm 0.022$ $0.290\pm 0.023$
Nash-MTL $\bm{0.970\pm 0.015}$
CAGrad $0.950\pm 0.022$ $0.689\pm 0.031$
1-step DiBS-MTL $0.890\pm 0.023$ $\bm{0.850\pm 0.022}$ $\bm{0.920\pm 0.025}$ $\bm{0.700\pm 0.037}$
1-step  DiBS-MTL    FAMO    UW    LS    FairGrad     Nash-MTL     CAGrad
(a) No task transformed
(b) reach transformed
(c) peg insert transformed
(d) window open transformed
Figure 3: Average task success rate on the MT10 benchmark, averaged over all tasks and 10 random seeds. Transparent curves show per-episode success rates, while solid curves denote rolling averages computed over a window of 51 episodes. 1-step DiBS-MTL demonstrates strong robustness to diverse reward transformations and consistently and significantly outperforms all baselines across all transformed settings.
5.2 Does the use of normalized gradients in DiBS-MTL induce robustness to nonaffine transformations?
We now conduct experiments in the multi-task reinforcement learning (MTRL) Meta-World MT10 benchmark (Yu et al., 2020b) to investigate how DiBS-MTL and baseline MTL methods perform when some tasks undergo nonaffine transformations. We limit our study in this section to MTRL, because task-specific losses in supervised learning examples such as computer vision are relatively standardized in the literature and thus transformations there will be unwarranted (Wang et al., 2022; Azad et al., 2023). In comparison, as we mention in Section 1, in the MTRL setting, rewards for different tasks might be designed on very different scales.
Metaworld MT10 (Yu et al., 2020b).
This benchmark consists of a robot arm conducting 10 tasks in the MuJoCo simulator (Todorov et al., 2012), each with distinct reward functions. The list of the 10 tasks is given in Appendix E.
MTRL baselines and metrics.
As all DiBS-MTL variants are transformation invariant, we restrict our study to 1-step DiBS-MTL, which has the lowest per epoch complexity. We compare 1-step DiBS-MTL with (1) LS, (2) UW, (3) FAMO, (4) Fairgrad (Ban and Ji, 2024), (5) Nash-MTL (Navon et al., 2022) and (6) CAGrad (Liu et al., 2021a). Following prior work, for every method, we use Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as the underlying reinforcement learning algorithm, hyperparameters values are taken from the corresponding baseline works, and we report the fraction of tasks successfully completed  (Navon et al., 2022; Liu et al., 2023), where the definition of task success is as described in the Meta-World benchmark (Yu et al., 2020b). Reported results have been averaged over 10 random seeds. Complete implementation details and hyperparameter values are in Appendix E.
Task transformations in the MTRL setup.
Reward transformations.
We consider the following task-transformation pairs: (i) $h(r)=\operatorname{sign}(r)\cdot r^{4}$ for reach, (ii) $h(r)=(5+r)^{4}$ for peg insert, ($h$ is monotonic over the reward range), and (iii) $h(r)=\exp(r)$ for window open. These transformations were arbitrarily selected to illustrate a range of functional forms and tasks.
DiBS-MTL displays robustness against nonaffine reward transformations, while baselines drastically degrade.
Figures 3(a), 3(b), 3(c) and 3(d) and Table 1 show that 1-step DiBS-MTL achieves remarkable robustness to the different reward transformations compared to existing MTL methods. While gradient-based MTL methods like FairGrad, Nash-MTL and CAGrad display strong performance in the nominal setting, their performance significantly degrades in the transformed cases, often to the extent that training became unstable, and the underlying MuJoCo simulator (Todorov et al., 2012) consistently gave out of bound errors for many seeds. Out of all methods, 1-step DiBS-MTL achieves state-of-the-art overall robustness for all transformed cases, while other methods either fail or exhibit significant performance drops for all transformations compared to their performance in the nominal setting. Finally, DiBS-MTL again achieves faster runtimes than the previous bargaining-inspired MTL method Nash-MTL (see Section F.1).
Takeaway from Sections 5.1 and 5.2: DiBS-MTL significantly outperforms state-of-the-art MTL baselines in all examples where underlying task losses/rewards are poorly scaled relative to each other, thanks to its monotonic transformation invariance. This highlights the unique suitability of DiBS-MTL to MTL problems with heterogeneous, possibly unknown task loss scalings.
Table 2: DiBS-MTL variants exhibit strong performance on the NYU-v2 benchmark, achieving state of the art values on MTL relevant metrics ($\Delta m\%$ & MR). Bold/underline indicates best/second-best performance per column. Results averaged over 3 random seeds.
@@ 5.2 Does the use of normalized gradients in DiBS-MTL induce robustness to nonaffine transformations?
STL 38.30 63.76 0.68 0.28 19.21 25.01 69.15 57.20 30.14
MGDA 32.03 60.77 0.6102 0.2453 19.00 24.64 69.83 57.78 30.55 -0.69 11.0
UW 39.08 64.73 0.5464 0.2284 23.04 27.34 62.85 49.23 23.49 3.77 8.2
Nash-MTL 40.16 65.65 0.5331 0.2203 19.96 25.25 68.29 55.72 28.62 -3.98 4.2
FAMO 37.58 64.08 0.5595 0.2296 19.15 25.04 69.36 57.44 30.23 -3.82 8.5
LS 40.16 65.63 0.5445 0.2223 23.03 27.50 62.66 49.32 23.61 3.06 6.0
GO4Align 38.87 64.52 0.5698 0.2406 23.41 27.81 61.83 48.56 23.45 5.47 9.9
IGBv2 33.61 61.88 0.5728 0.2222 23.18 27.94 62.14 49.10 23.59 11.92 9.7
IGBv2+Nash 38.81 65.66 0.5258 0.2086 20.09 25.50 67.82 55.35 28.37 1.67 5.1
LDC 39.21 65.70 0.5337 0.2180 20.23 25.29 67.98 55.19 27.98 -3.24 4.5
1-step DiBS-MTL 40.92 66.60 0.5337 0.2216 20.06 25.35 68.15 55.54 28.54 -4.11 3.1
5-step DiBS-MTL 41.27 67.10 0.5473 0.2300 19.69 25.13 68.69 56.28 29.31 -4.57 4.2
10-step DiBS-MTL 41.33 67.09 0.5381 0.2260 19.77 25.06 68.67 56.12 29.02 -4.74 3.6
Table 3: DiBS-MTL exhibits strong performance on the QM9 benchmark, with the 10-step variant achieving best and second-best values on MTL relevant metrics $\Delta m\%$ & MR. Bold/underline indicates best/second-best performance. Results averaged over 3 random seeds.
$\mu$ $\alpha$ $\epsilon_{\text{HOMO}}$ $\epsilon_{\text{LUMO}}$ $\langle R^{2}\rangle$ ZPVE $_{0} UU$ $H$ $G$ $c_{v}$ $\Delta m\%\downarrow$ MR $\downarrow$
MAE $\downarrow$
STL 0.07 0.18 60.60 53.90 0.50 4.53 58.80 64.20 63.80 66.20 0.07
LS 0.0998 0.3108 69.83 84.87 5.13 13.43 137.25 137.97 138.43 133.55 0.1226 167.16 7.42
GO4ALIGN 0.1047 0.3175 70.40 86.48 5.05 13.35 138.78 139.49 139.86 135.27 0.1252 167.90 7.88
UW 0.4123 0.3937 159.07 145.41 1.09 4.85 61.52 61.95 61.98 61.40 0.1206 102.96 5.50
LDC 0.2782 0.3318 141.53 121.37 1.20 4.34 51.43 51.83 51.91 51.28 0.1076 68.79 3.50
FAMO 0.4422 0.3998 158.41 145.83 1.10 5.04 65.76 66.15 66.24 65.40 0.1267 110.99 6.67
Nash-MTL 0.1024 0.3153 70.75 85.29 5.05 13.04 135.67 136.55 136.82 132.71 0.1227 165.12 6.96
IGBV2 0.2866 0.3557 149.29 135.04 1.24 4.60 64.68 65.08 65.15 64.92 0.1176 85.75 5.33
IGBV2 + Nash 0.4399 0.3762 142.19 139.73 1.62 5.10 67.20 67.61 67.66 67.13 0.1159 115.24 7.17
1-step DiBS-MTL 0.1347 0.2841 97.53 95.01 2.56 5.75 65.92 66.38 66.39 65.83 0.1067 71.83 4.83
5-step DiBS-MTL 0.1285 0.2914 96.40 95.95 2.55 5.68 69.28 69.68 69.74 69.11 0.1065 72.92 5.54
10-step DiBS-MTL 0.1213 0.2860 93.50 90.46 2.37 5.81 70.12 70.54 70.75 69.83 0.1061 67.37 5.21
5.3 Can the transformation robust DiBS-MTL maintain competitive performance in untransformed settings on standard MTL benchmarks?
@@
To compare DiBS-MTL with existing methods in large-scale learning examples, we perform experiments in supervised learning in two domains—(i) computer vision (the NYU-v2 dataset (Silberman et al., 2012)), and (ii) graph neural networks for quantum chemistry (QM9 dataset (Blum and Reymond, 2009)). Additional results for the computer vision Cityscapes dataset (Cordts et al., 2016) are included in Appendix D. All of these are widely used challenging benchmarks in MTL literature (Liu et al., 2021a; Navon et al., 2022; Liu et al., 2023; Xiao et al., 2025; Dai et al., 2023; Shen et al., 2024; Ban and Ji, 2024), and complete benchmark descriptions are given in Appendix E. We begin by describing the baselines and metrics for all datasets. We keep the default training procedures established in prior baseline works for a fair comparison. All implementation details are given in Appendix E.
Supervised MTL baselines.
For all datasets, we compare 1-Step, 5-Step, and 10-Step DiBS-MTL with an extensive suite of strong MTL baselines: (1) linear scalarization (LS), (2) uncertainty weighting (UW) (Kendall et al., 2018), (3) MGDA (Sener and Koltun, 2018), (4) Nash-MTL (Navon et al., 2022), (5) FAMO (Liu et al., 2023), (6) GO4Align (Shen et al., 2024), (7) IGBv2, (8) IGBv2+Nash (Dai et al., 2023), and (9) LDC (Xiao et al., 2025). For all baselines across all datasets, the architecture, data pipeline, protocols, and training hyperparameters are taken to be the same as that in the corresponding works introducing the baselines.
Supervised MTL metrics.
In addition to task-specific metrics, we evaluate all methods using two standard measures of multi-task performance: (1) the mean relative per-task performance drop $\Delta m\%$ with respect to single-task learning (STL), and (2) the mean rank (MR) across all task metrics. $\Delta m\%$ captures the average improvement or degradation relative to independent single task training (STL), while MR summarizes the average rank of a method across all tasks for a benchmark. These metrics are used in all baseline works as well, to assess whether an MTL method improves joint learning across diverse tasks.
DiBS-MTL achieves competitive performance across all datasets.
Table 2 reports results on NYU-v2. DiBS-MTL variants achieve state-of-the-art performance in $\Delta m\%$ as well as MR, being the best or second-best in both MTL relevant metrics across all methods, demonstrating balanced improvements across segmentation, depth, and normal prediction. DiBS-MTL variants also achieve competitive or superior performance in segmentation and depth estimation, while matching strong baselines on surface normal prediction. In the QM9 benchmark, DiBS-MTL variants display superior performance for the $\alpha$ and $c_{v}$ prediction tasks. Additionally, DiBS-MTL variants achieve the best $\Delta m\%$ and second best MR performance (Table 3). Finally, 1-step DiBS-MTL is faster than Nash-MTL for NYU-v2, and only slightly slower on QM9 (Section F.1).
Takeaway from Section 5.3: Across all large-scale untransformed benchmark evaluations, DiBS-MTL  consistently performs competitively against leading MTL baselines. This indicates that the transformation robustness property of DiBS-MTL does not sacrifice performance on popular real-world benchmarks.